a class of 30 of these pupils is used as sample. what is the probability that the class mean is greater than 46 inches?


1
Expert's answer
2021-03-04T23:50:16-0500

Given,

Mean "\mu=45"

Standard deviation "\sigma=2"


Probability that the class mean is greater than 46 ="P(X>46)"


"z=\dfrac{x-\mu}{\sigma}=\dfrac{46-45}{2}=\dfrac{1}{2}=0.5"



Therefore the required probability is "P(z>0.5)=P(0<z<\infty)-P(0<z<0.5)"

"=0.5-0.1915=0.3085"